OpenStudy (anonymous):

The data 6, 12, x, and 7 have a mean of ten. Find the value of x.

OpenStudy (ghazi):

mean= (6+12+x+7)/4=10 now solve for x

OpenStudy (anonymous):

i got 6.25 is that right?

OpenStudy (ghazi):

nope...one more hint 6+12+x+7=10*4=40

OpenStudy (ghazi):

25+x=40 now solve for x

OpenStudy (anonymous):

okay i got 15

OpenStudy (ghazi):

you're done :)

OpenStudy (anonymous):

thanks!!

OpenStudy (ghazi):

anytime